OpenStudy (anonymous):

Cot x sec^4x = cot x + 2 tan + tan^3

OpenStudy (anonymous):

And?

OpenStudy (anonymous):

its a true statement ; note that\[\sec^4 x=\sec^2 x \ \sec^2 x=(1+\tan^2 x)^2\]

OpenStudy (anonymous):

I need help figuring it out, i spent five hours and still did not come up with an answer

OpenStudy (anonymous):

Using mukushla hint, it is simple. 1/Tanx * (1+tan^x)^2 = 1/Tanx + 2 tanx + tan^3x and cross multiply (1+tan^x)^2 = 1 + 2 Tan^2 x + tan^4x QED